# Rui Eduardo Lopes, Student Member, IEEE, Susana Sargento, Senior Member, IEEE

### v1

@@ II-C Industrial landscape
These individual initiatives are being target towards a cooperation. By 2024, at CES 2024, AUTOSAR, Connected Vehicle Systems Alliance (COVESA), Eclipse SDV, and SOAFEE announced a “collaboration of collaborations” to “align efforts in the SDV ecosystem and create a joint SDV vision clear and unified definition of what constitutes an SDV, which points towards a brighter future in SDV with greater interoperability [35].
@@
III Main building blocks of Software Defined Vehicles
@@
Developing a vehicle within the SDV paradigm ( a ) requires considerations on several verticals within the context provided by fig. 5:
@@
  • Service Oriented Architectures ( b section III-A): how vehicular services are developed and communicate between one another and abstractions for the vehicular service development and for transferring data between services.
@@
  • Virtualization techniques and operating systems ( c , d sections III-B and III-C): how vehicular services are deployed within each ECU, including the software stack within each ECU and virtualization techniques..
@@
  • Intra and inter vehicular communication architectures, protocols and physical layers ( e section III-D): how vehicular ECUs communicate with each other — intra-vehicular communications — and with ECUs on other vehicles — inter-vehicular and deterministic communications ( f section III-E): how to ensure those communications always happen and within the necessary reliability and timing requirements.
Existing automakers were not agile enough within the development of their vehicular services [37], which lead to the rise of new automakers, such as Tesla, in very software-intensive tasks such as autonomous driving.
The SDV paradigm main ideas are not new. Prior to the advent of SDVs, the automotive industry focus on software was mostly restricted to connectivity between vehicles and the user’s smartphones to provide basic insights into the state of the vehicle [38]. By 1998, [39] considered intelligent vehicles as essential to autonomous transportation, designing a proposal of a vehicle capable of communicating with a centralized management control system. Also in 1998, [40] proposes that smart vehicles are a junction of communication architectures, system architecture, sensors, tasks, kinematics and man-machine interface. The authors propose a communication architecture in a producer-distributor-consumer pattern with buses such as RS232 and CAN, where time coherence is necessary. By 2001, a reference model architecture for intelligent vehicles [41] considered the vehicle as having intelligent control systems with each layer with one or more computational nodes interconnected and connected to sensors.
III-A Service-Oriented Architecture
With vehicular E/E architectures getting more complex, standards were required in order to model both the ECUs hardware and the software. The AUTOSAR is a initiative of vehicle manufacturers based on design by contract [8, 42] for the development and modeling of vehicular applications through Object Oriented Programming (OOP) concepts. During the 2000s this initiative aimed at defining the whole vehicular software stack with support to running several applications and communication with other services through several buses — either CAN, LIN or even Ethernet — through an abstraction of the several hardware possibilities into a Run Time Environment (RTE[8].
While a very mature de facto standard with the support from vehicle manufacturers and vendors, AUTOSAR is a highly complex ecosystem, with more than 100 modules and its documentation reaching more than 21 000 pages [8]. Development through the original Classic Platform is also very hard-coded to the vehicle and its E/E architecture [43]. AUTOSAR has two different types of platforms as depicted in fig. 6: the already mentioned Classic, for traditional hard real-time systems based on micro-controllers with a static configuration, and set of pre-compiled monolithic binaries; but also a newer Adaptive platform, for high performance services and ECUs, with greater flexibility and adaptability to new or changed services. The Classic Platform is aimed at safety critical systems with hard real-time requirements with a stated linkage between services and CPU cores, while the Adaptive is aimed at high performance ECUs [14] and is capable of considering heterogeneous ECUs and departs from legacy buses like CAN towards a SOA communication architecture with Ethernet based on Scalable service-Oriented MiddlewarE over IP (SOME/IP) as bus protocol [44] and is overall more dynamic but less time predictable [45, 29], essential within critical vehicular systems.
Figure 6: AUTOSAR Classic VS Adaptive.
The SOAFEE [^22] Available at https://www.soafee.io/. More details at https://www.arm.com/blogs/blueprint/cloud-native-automotive-development presents a cloud-native alternative with an architecture enhanced for mixed-criticality automotive applications with corresponding open-source reference implementations to enable commercial and non-commercial offerings.
Each ECU can now part of a SOA, where distributed services interact with one another based on some remote access protocol. In this context, and as defined by [46], a service is a mechanism to enable access to one or more capabilities, where the access is provided using a prescribed interface and is exercised consistent with constraints and policies as specified by the service description. Figure 7 presents an overview of the SOA paradigm.
@@ III-A Service-Oriented Architecture
Figure 7: SOA as perceived by the Renault Group automaker (adapted). Services are the basic unit of vehicular applications, from simple services that directly interact with sensors and actuators, up to more complex services — deployed either in vehicle or in a cloud — that integrate simpler service. Services communicate with one another via one or more message bus [19].
Service Oriented Architectures are a software architectural pattern where each component provide and consume services — a discrete unit of functionality — over a predefined protocol over a network. It can be considered an architecture following the main definitions of the Actor Model computational model, since components are decoupled from how the communication is achieved, allowing asynchronous communication [47]. SOA, when implemented on a vehicle, on a more reduced set of zonal ECUs, bring benefits such as the decoupling of hardware and software, increased modularity and separation of concerns, and contribute towards the overall decrease in hardware complexity. [48]:
@@
At the cost of increased complexity in maintaining contracts between services and ensuring availability and discovery of the services [48]. This is particularly complex within a vehicular context, with a SOA supporting at most 60 to 80 services [24], not addressing situations where there are hundreds or thousands of micro-services.
Within SOA based architectures, data layers can abstract the physical medium of services by providing a unified way of exchanging and accessing data between services on different components/ECUs [49]. Three main design paradigms can be considered for implementing a data layer [50] a) a fully periodic publisher-subscriber based, b) a fully event triggered client-server based, or c) a mixed model, with periodic model for non time-critical chains and event-triggered for time-critical chains.
Information from vehicle sensors can be fed from/to services leveraging the ECUs and their communication paradigm through data layer solutions like DDS [^23] Available at https://www.dds-foundation.org/. middleware, SOME/IP, Constrained Application Protocol (CoAP) or Message Queuing Telemetry Transport (MQTT[29].
@@
Publish/subscribe paradigms — with protocols such as MQTT, DDS or SOME/IP — can respond to a data centric paradigm, when the data sending is event triggered, unlike solutions such as Hypertext Transfer Protocol (HTTP) or CoAP based on the request/response paradigm [25].
Table II provides a summary on some of the main possible solutions for SOA data middleware.
TABLE II: Qualitative comparison of IoT standards (partially based on [51, 52, 53]).
@@
[t] Standard MQTT DDS SOME/IP Zenoh HTTP Transport UDP (in MQTT-SN), TCP UDP, TCP IP UDP, TCP, QUIC TCP, QUIC Paradigm Pub / Sub Pub / Sub Pub / Sub Pub / Sub Data storage Request / Reply Node Discovery No Yes (SDP and alternatives) Yes Yes No Fault Tolerance Broker is SPoF Yes, fully distributed by default Yes Yes, can be fully distributed. Client and server are SPoF Data Schema None Required Required Only encoding is required Optionalc QoS Policies At delivery level At multiple entitiesa N/A At 2 levelsb N/A
@@
While all these solutions can ensure delivery of data, there is a tendency to consider DDS and Zenoh the best performant solutions since they provide better flexibility and performance. According to [54], which compared performance of MQTT, DDS, and Zenoh, the latter shows (for multi-machine scenarios, more realistic in zonal vehicular architectures) a double increase on the throughput of DDS and 50 times of MQTT, while DDS usually provides the best latency and better flexibility with more QoS levels.
III-B Virtualization and orchestration techniques
@@
Virtualization provides support for running multiple OSs on the same hardware. It is an essential concept in the context of vehicular systems, since in zonal architectures the number of ECUs will tend to be much less than the number of services and required OSs.
Virtualization allows hardware to be shared and not uniquely assigned to one operating system, while achieving the necessary isolation automakers require between critical and non-critical services. Within the responsibilities of the virtualization layer is the abstraction of hardware and of network topologies [55]. Virtualization also allows to deploy services in heterogeneous platforms and ECUs [56].
@@
  • Shared I/O: how to distribute input and output devices, such as Universal Serial Bus (USB) or Ethernet buses, across virtual machines. Allowing virtualized services to use and share the same Ethernet port for example requires a scheduler that is capable of communicating with the different virtual machines and the deterministic network configurator. This communication can be achieved through hypercalls, the equivalent of syscalls but to invoke a hypervisor [57]. This may impact however the overall frame transmission delays.
@@
Vehicular OSs such as Android Automotive provides full support for virtualization through Kernel-based Virtual Machine (KVM) and virtIO [^25] Developer documentation available at https://wiki.libvirt.org/Virtio.html devices, allowing support at the Virtual Machine (VM) side for sensors, graphics, storage, network (including Wi-Fi through virtio-net and VirtWifi [^26] Developer documentation available at https://source.android.com/docs/devices/automotive/ virtualization/architecture.) and with support for communicating with the vehicle through Google RPC (gRPC) [^27] Details available at ..
@@
Eclipse, within the SDV foundation proposes Eclipse Ankaios [^28] Developer documentation available at https://projects.eclipse.org/projects/automotive.ankaios to “manage multiple nodes and virtual machines with a single unique API in order to start, stop, configure, and update containers and workloads”. Ankaios is tailored towards automotive applications and developed for automotive ECUs and provides a central place to manage automotive applications, supporting:
@@
Kubernetes [^29] Available at https://kubernetes.io/., as a container orchestration tool mainly designed for datacenters, can also be consider to manage vehicular workloads. The work of [58] proposes vehicles as nodes of a Kubernetes cluster.
@@
Apart from its inherent advantages as container orchestration and scheduler, Kubernetes is very flexible in networking terms. Networking between containers can be changed through Container Network Interfaces [^30] Project available at https://github.com/containernetworking/cni., a specification for configuring network interfaces in containers. This allows considering deterministic networking concepts — further discussed in section III-E — in container scenarios. For example the work of [59] proposes support for deterministic networking within Kubernetes containers without modifying the application binaries — essential to ensure compatibility with existing vehicle services but also to simplify the development process of new services — through a side car container that exposes the standard Portable Operating System Interface (POSIX) socket interface to other containers and redirects traffic to deterministic configured network interfaces in the host computer. Integration with cloud infrastructure is also possible. For example, within [60] is proposed a full-fledged dynamic interference-aware cloud scheduling architecture for latency-sensitive workloads.
III-C Vehicular operating systems
ECUs within a vehicle run services that require deterministic scheduling from the OS. Failure to have a deterministic real-time scheduling can lead to services being dispatched too soon or too late to the OS, with dramatic effect on the services behavior. Therefore, Real-Time Operating System are a need within vehicular computing domains to ensure services are dispatched in a determined specific order and time-frame. While many OSs can support real-time constraints — including any generic Linux kernel based OS with PREEMPT_RT [^31] Available at https://wiki.linuxfoundation.org/realtime/. patches [61], some solutions are heavily automotive focused, such as Android Automotive, Nvidia Drive OS and QNX.
Android Automotive [^32] Developer documentation available at https://source.android.com/docs/devices/automotive. Not to be confused with Android Auto, a platform to use a smartphone as a infotainment alternative. and its extensions such as SnapOS [^33] Available at https://www.snappautomotive.io/snappos provide an OS and a platform that is based on Android — and can therefore leverage the existing application ecosystem — while supporting specific characteristics of vehicles, including the support for the creation of a vehicle abstraction layer required to abstract the vehicle’s specific/proprietary components, UI support for setting vehicle’s settings and OTA updates.
Nvidia Drive OS [^34] Developer documentation available at https://developer.nvidia.com/drive/driveos is the “reference operating system and associated software stack designed specifically for developing and deploying autonomous vehicle applications on DRIVE AGX-based hardware”. It is in fact a custom RTOS with a Nvidia Hypervisor and necessary CUDA and TensorRT libraries to fully utilize the hardware (graphical) potential.
Yocto [^35] Available at https://www.yoctoproject.org/-based OSs are also a possibility. Yocto supports the creation of custom OS based on Linux, allowing adding custom layers to embedded systems. Yocto based Wind River Linux [^36] Available at https://www.windriver.com/products/linux. is an example of such family of OS, providing support for multiple board support packages in both x86 and Arm hardware. Automotive Grade Linux [^37] Available at https://docs.automotivelinux.org/, https://github.com/nxp-auto-linux/auto_yocto_bsp. is another example of a Yocto based OS and a industry supported initiative to develop software for the “infotainment, instrument cluster, Heads-Up Display (HUD), telematics, connected car, ADAS, functional safety, and autonomous driving” [62].
Alternative OSs include the VxWorks [^38] Available at https://www.windriver.com/products/vxworks [63], a RTOS for mission critical embedded systems. While usually considered in spacial scenarios, automotive are also supported for example through the integration with AUTOSAR Adaptive Platform. The QNX [^39] Available at https://blackberry.qnx.com. is another alternative as a micro-kernel RTOS with extensive foothold in the automotive sector installed in more than 215 million vehicles. Another open source initiative, albeit in initial stages (launched by January 2024), is the Eclipse ThreadX [^40] Previously Azure RTOS. Available at https://threadx.io/, with developer documentation at https://projects.eclipse.org/proposals/eclipse-threadx., an open-source, vendor neutral RTOS.
Vehicular services should be developed independently of the RTOS and thus an isolation between the services and the RTOS is desired. Initiatives such as Operating system abstraction layer (OSAL) from IBM [^41] Developer documentation at https://www.ibm.com/docs/en/engineering-lifecycle-management-suite/design-rhapsody/9.0.2?topic=references-operating-system-abstraction-layer-osal. or NASA consider “isolates embedded application software from a RTOS by providing “a well-defined, generic interface to RTOS services” [^42] More information at https://opensource.gsfc.nasa.gov/projects/osal/. and can, therefore, contribute to the necessary portability of services between RTOS.
These operating systems, being mostly RTOS or RTOS-like, can be evaluated through the worst-case execution time for each task, modularity, response time or jitter [64]. At the time of writing, no comparison between some or all of the aforementioned OSs was found, with the choice of OS mostly related to the associated ecosystem (such as already available applications or SDKs).
III-D Inter- and intra-vehicular communication

### v2

@@ 2.1 Industrial landscape
These individual initiatives are being target towards a cooperation. By 2024, at CES 2024, AUTOSAR, Connected Vehicle Systems Alliance (COVESA), Eclipse SDV, and SOAFEE announced a “collaboration of collaborations” to “align efforts in the SDV ecosystem and create a joint SDV vision clear and unified definition of what constitutes an SDV, which points towards a brighter future in SDV with greater interoperability [33]. Innovation projects such as the HAL4SDV212121hal4sdv.eu also aim to contribute to a more holistic solution by including in its goals a ”standardized interfaces to sensors, actors, compute resources and persistent storage”.
@@
3 Main building blocks for Software Defined Vehicles
@@
Developing a vehicle within the SDV paradigm ( a ) requires considerations on several domains within the context provided by fig. 7, including:
  • 1.
    Physical disposition of the vehicular ECUs and specifications of each ECU ( c section 3.1): this domain explores how ECUs are physically disposed throughout the vehicle, namely considering a zonal Electrical/Electronic architecture.
  • 2.
@@
    Service Oriented Architectures ( b section 3.2): this domain explores how vehicular services are developed and communicate between one another and abstractions for the vehicular service development and for transferring data between services.
  • 3.
@@
    Virtualization techniques and operating systems ( d , e sections 3.4 and 3.3): this domain considers how vehicular services are deployed within each ECU, including the software stack within each ECU and virtualization techniques.
  • 4.
@@
    Intra and inter vehicular communication architectures, protocols and physical layers ( f sections 3.5 and 3.6 this domain considers how vehicular ECUs communicate with each other and with ECUs on other vehicles (intra-vehicular and inter-vehicular communications, respectively).
  • 5.
    Deterministic vehicular communications ( g section 3.7): this domain considers how to ensure vehicular communications always happen and within the necessary reliability and timing requirements.
The next subsections further detail each one of these building blocks, including their state of the art and associated main challenges.
3.1 Physical E/E architectures
Figure 8: Legacy distributed, domain, and zonal E/E architectures within a vehicle.
To support SDVs, the Electrical/Electronic and computational architectures within vehicles must also evolve from the fully distributed vision, with ECUs scattered throughout the vehicle towards a more centralized solution, with zone-based architectures, i.e. ECUs positioned in specific parts of the vehicle responsible for certain types of actions [35]. Figure 8 provides an overview of this evolution.
During the 1980s (fig. 8a)), new features were added at the cost of adding new specialized ECUs , close to the sensors and actuators and connected through CAN or other equivalent bus. With the explosion of features in vehicles, especially luxury cars, during the 1990s, the number of ECUs increased significantly, making vehicles more complex, and thus increasing the amount of systems that need to be choreographed, developed, tested and that have to communicate, as well as the possibilities of something going wrong [36, 37].
By the 2000s, the high complexity of intra vehicular systems became subject to de facto standards such as AUTOSAR [36, 37] and buses such as FlexRay aimed at mitigating some of the shortcomings of CAN. Since the ECUs were scattered throughout the vehicle, with little to no centralization, these legacy architectures were distributed, allowing a better decoupling, separation of concerns and higher tolerance to equipment failure. However, challenges such as the scalability of the deploying services across multiple ECUs (considering hardware variations and the absense of virtualization), communication performance, and the high costs of the associated complex, bulky, wiring harnesses have driven a trend towards a more centralized architecture, consolidating processing into fewer computing units.
A centralized architecture can reduce costs and complexity by decreasing the number of ECUs, increasing software reusability, and enabling greater flexibility.The centralized E/E architectures supporting SDVs follow the principle of aggregation of ECUs into a small subset of more powerful High Performance Computers. Each HPC can potentially contain multi core Central Processing Units, Graphical Processing Units, Digital Signal Processors, and potentially Field-Programmable Gate Arrays [36, 37, 38], as well as high performance network capabilities (e.g. Ethernet). Some ECUs can eventually be dedicated to computing intensive tasks, including Machine Learning (ML) and Deep Learning (DL) workloads supporting ADAS systems [39]. This follows previous trends within other industries, in particular the aviation industry [19].
Centralization within the intra-vehicular systems can be achieved at a domain, cross domain or zone/vehicle centralized way [19]. By 2010s, domain and cross domain architectures (fig. 8b)) were a first step towards mitigating those issues through the aggregation of functions per domain. In a domain-oriented solution, the grouping of ECUs into HPCs is done by vehicle function since each function requires heavy interaction between ECUs (specially in complex systems such as ADAS). In a cross-domain solution, functions of several domains are consolidated within more powerful ECUs. This minimizes, without removing, the high bandwidth required for communication between ECUs. However, the integration of multiple domains required by systems such as ADAS still require coordination between different domains [36, 37].
Zone-oriented architectures (fig. 8 (b)) aimed to solve the issue of coordination between domains. Each zone of the vehicle collects and forwards information to a much more powerful central ECU, or even offloads to the cloud. Bosch reported on achieving a 15–20% reduction in wiring harness weight when using a zone-oriented architecture instead of a domain-oriented architecture, showcasing the potential of zonal architectures versus the previous approaches.
Future approaches should also consider other vehicles and infrastructure (e.g. smart cities) as an integral part of the SDV environment services can consider for deployment.The next phases of automotive software development should also have a deep integration with cloud and follow the best practices of software development (such as CI/CD) [40] and should consider deterministic networking concepts and standards, as discussed in section 3.7.
3.2 Service-oriented architectures
As vehicular E/E architectures become more complex, standards and well defined patterns are required in order to model both the ECUs hardware and the software. Service Oriented Architectures are an example of such architectural pattern, where each component provide and consume services. Figure 9 presents an overview of the SOA paradigm. Services in this context are a discrete unit of functionality, from directly interacting with sensors and actuators, up to more complex interactions between simpler services. A service contains the mechanism to enable access to one or more capabilities, where the access is provided using a prescribed interface and is exercised consistent with constraints and policies as specified by the service description [41].
@@ 3.2 Service-oriented architectures
Figure 9: SOA in vehicles as perceived by the Renault Group automaker (adapted). Services (in red) are the basic unit of vehicular applications and communicate with one another via one or more message bus [14].
@@
When implemented on a vehicle, on a more reduced set of (zonal) ECUs, the SOA paradigm brings benefits such as the decoupling of hardware and software, increased modularity and separation of concerns, and contribute towards the overall decrease in hardware complexity. [42]. These advantages are, however, at the cost of increased complexity in maintaining contracts between services and ensuring availability and discovery of the services [42]. This is particularly complex within a vehicular context, with a SOA supporting at most 60 to 80 services [23], not addressing situations where there are hundreds or thousands of micro-services.
Within SOA based architectures, data layers (message bus in fig. 9 can abstract the physical medium of services by providing a unified way of exchanging and accessing data between services on different components/ECUs [43]. Information from vehicle sensors can then be fed from/to services leveraging the ECUs and their communication paradigm through data layer solutions.
@@
Three main paradigms can be considered for implementing a data layer [44]: a) a fully periodic publisher-subscriber based or b) a fully event triggered client-server based. Examples for the publish/subscribe paradigms include the DDS [^22] At dds-foundation.org. middleware, SOME/IP, or Message Queuing Telemetry Transport (MQTT[22]. Publish/subscribe paradigms can respond to a data centric paradigm, when the data sending is event triggered, unlike solutions such as Hypertext Transfer Protocol (HTTP) or Constrained Application Protocol (CoAP) based on the request/response paradigm [24]. Abstraction layers such as uprotocol [^23] At projects.eclipse.org/projects/automotive.uprotocol. can be considered to decouple vehicular services from the chosen data layer.
Table 2 provides a summary on some of the main possible solutions for SOA data middleware. While all these solutions can ensure delivery of data, there is a tendency to consider DDS and Zenoh the best performant solutions since they provide better flexibility and performance. According to [45], which compared performance of MQTT, DDS, and Zenoh, the latter shows (for multi-machine scenarios, more realistic in zonal vehicular architectures) a double increase on the throughput of DDS and 50 times of MQTT, while DDS usually provides the best latency and better flexibility with more Quality of Service (QoS) levels.
Table 2: Qualitative comparison of IoT standards (partially based on [46, 47, 48]).
@@
[t] Standard MQTT DDS SOME/IP Zenoh HTTP Transport UDP (in MQTT-SN), TCP UDP, TCP IP UDP, TCP, QUIC TCP, QUIC Paradigm Pub / Sub Pub / Sub Pub / Sub Pub / Sub Data storage Request / Reply Node Discovery No Yes (SDP and alternatives) Yes Yes No Fault Tolerance Broker is SPoF Yes, fully distributed by default Yes Yes, can be fully distributed Client and server are SPoF Data Schema None Required Required Only encoding is required Optionalc QoS Policies At delivery level At multiple entitiesa N/A At 2 levelsb N/A
@@
3.3 Vehicular operating systems
ECUs within a vehicle run services that require deterministic scheduling from the OS. Failure to have a deterministic real-time scheduling can lead to services being dispatched too soon or too late to the OS, with dramatic effect on the services behavior. Therefore, Real-Time Operating System are a need within vehicular computing domains to ensure services are dispatched in a determined specific order and time-frame. While many OSs can support real-time constraints — including any generic Linux kernel based OS with PREEMPT_RT [^24] Available at wiki.linuxfoundation.org/realtime. patches [49], some solutions are heavily automotive focused, such as Android Automotive, Nvidia Drive OS and QNX.
Android Automotive [^25] Not to be confused with Android Auto, a platform to use a smartphone as a infotainment alternative. and its extensions such as SnapOS [^26] Available at snappautomotive.io. provide an OS and a platform that is based on Android — and can therefore leverage the existing application ecosystem — while supporting specific characteristics of vehicles, including the support for the creation of a vehicle abstraction layer required to abstract the vehicle’s specific/proprietary components, UI support for setting vehicle’s settings and OTA updates.
Nvidia Drive OS [^27] Dev documentation at developer.nvidia.com/drive. is the “reference operating system and associated software stack designed specifically for developing and deploying autonomous vehicle applications on DRIVE AGX-based hardware”. It is in fact a custom RTOS with a Nvidia Hypervisor and necessary CUDA and TensorRT libraries to fully utilize the hardware (graphical) potential.
Yocto [^28] Available at yoctoproject.org. based OSs are also a possibility. Yocto supports the creation of custom OS based on Linux, allowing adding custom layers to embedded systems. Yocto based Wind River Linux [^29] At windriver.com. is an example of such family of OS, providing support for multiple board support packages in both x86 and Arm hardware. Automotive Grade Linux [^30] At docs.automotivelinux.org, github.com/nxp-auto-linux. is another example of a Yocto based OS and a industry supported initiative to develop software for the “infotainment, instrument cluster, Heads-Up Display (HUD), telematics, connected car, ADAS, functional safety, and autonomous driving” [50].
Alternative OSs include the VxWorks [^31] Available at windriver.com. [51], a RTOS for mission critical embedded systems. While usually considered in space scenarios, automotive are also supported for example through the integration with AUTOSAR Adaptive Platform. The QNX [^32] Available at blackberry.qnx.com. is another alternative as a micro-kernel RTOS with extensive foothold in the automotive sector installed in more than 215 million vehicles. Another open source initiative, albeit in initial stages (launched by January 2024), is the Eclipse ThreadX [^33] Previously Azure RTOS, now threadx.io, with dev documentation at eclipse.org., an open-source, vendor neutral RTOS.
These operating systems, being mostly RTOS or RTOS-like, can be evaluated through the worst-case execution time for each task, modularity, response time or jitter [52]. At the time of writing, no comparison between some or all of the aforementioned OSs was found.
Vehicular services should be developed independently of the RTOS and thus an isolation between both is desired. Initiatives such as the Operating system abstraction layer (OSAL) from IBM [^34] Dev documentation at ibm.com. or NASA “isolates embedded application software from a RTOS by providing “a well-defined, generic interface to RTOS services” [^35] More information at opensource.gsfc.nasa.gov. and can, therefore, contribute to the necessary portability of services between OSs. Without such initiatives, the definition of the OS will be mostly related to the associated ecosystem, hindering possiblity interoperability between services developed for each OS.
3.4 Virtualization and orchestration techniques
To ensure optimal usage of ECUs, with concurrent execution of different services with different criticality, partitioning of scheduling of each core per service(s) is required. Virtualization of resources, either memory, buses (e.g. CAN) and CPU allow both isolation of services within virtual machines while enabling migration and reusage of legacy software [19].
@@ 3.4 Virtualization and orchestration techniques
Virtualization provides support for running multiple OSs on the same hardware, ensuring the abstraction of hardware and of network topologies [53].One of the key advantages is allowing hardware to be shared and not uniquely assigned to one operating system, while achieving the necessary isolation automakers require between critical and non-critical services. It is an essential concept in the context of vehicular systems not only because of isolation but to support the aggregation of functions in less ECUs (per the zonal architectures) than the number of services and required OSs [54].
@@
  • 2.
    Shared I/O: how to distribute input and output devices, such as Universal Serial Bus (USB) or Ethernet buses, across virtual machines.
  • 3.
@@
Allowing virtualized services to use and share the same Ethernet port for example requires a scheduler that is capable of communicating with the different virtual machines and the deterministic network configurator. This communication can be achieved through hypercalls, the equivalent of syscalls but to invoke a hypervisor [55]. This may impact, however, the overall frame transmission delays.
@@
Vehicular OSs such as Android Automotive provides full support for virtualization through Kernel-based Virtual Machine (KVM) and virtIO [^37] Developer documentation at wiki.libvirt.org. devices, allowing support at the Virtual Machine (VM) side for sensors, graphics, storage, network (including Wi-Fi through virtio-net and VirtWifi [^38] Developer documentation at android.com.) and with support for communicating with the vehicle through Google RPC (gRPC) [^39] Details at android.com..
@@
Eclipse, within the SDV foundation proposes Eclipse Ankaios [^40] Developer documentation at projects.eclipse.org/projects/automotive.ankaios. to “manage multiple nodes and virtual machines with a single unique API in order to start, stop, configure, and update containers and workloads”. Ankaios is tailored towards automotive applications and developed for automotive ECUs and provides a central place to manage automotive applications, supporting:
@@
Kubernetes [^41] Available at kubernetes.io., as a container orchestration tool mainly designed for datacenters, can also be consider to manage vehicular workloads. The work of [56] proposes vehicles as nodes of a Kubernetes cluster.
@@
Apart from its inherent advantages as container orchestration and scheduler, Kubernetes is very flexible in networking terms, since networking between containers can be changed through Container Network Interfaces [^42] Project available at github.com/containernetworking/cni.. CNIs are a specification for configuring network interfaces in containers. This allows considering deterministic networking concepts (section 3.7) in container scenarios. For example, the work of [57] proposes support for deterministic networking within Kubernetes containers without modifying the application binaries through a side car container that exposes the standard Portable Operating System Interface (POSIX) socket interface to other containers and redirects traffic to deterministic configured network interfaces in the host computer. Not modifying the application binaries is essential to ensure compatibility with existing vehicle services but also to simplify the development process of new services.
3.5 Intra-vehicular communications